What value of b will make the expression x^2-bx+40 a perfect square?
Answer by stanbon(75887)

What value of b will make the expression x^2-bx+40 a perfect square?
------------
Perfect Square form: (c-d)^2 = c^2-2cd+d^2
Your Problem:
d^2 = 40
d = 2sqrt(10)
-----------------
c^2=x^2
c=x
==============
-2cd=-bx
Substitute to get:
bx = 2x(2sqrt(10))
bx = 4sqrt(10)x
b = 4sqrt(10)
=========================
Cheers,
Stan H.